The difference of two numbers is 5. The sum of the numbers is 35. Find the numbers using algebra.
Found 2 solutions by jorel1380, Love math:
Answer by jorel1380(3719)

x+5=y
x+y=35
x+x+5=35
2x=30
x=15
x+5=20..

Answer by Love math(28)

x-y =5
Add x+y =35
-------
2x = 40
x =40/2
x =20
plug x=20 solve for y
20+y =35
y= 35/20
y= 7/4